how do i write the linear equations in slope -intercept form for 16=0.5x-2y-2
Found 2 solutions by stanbon, oberobic:
Answer by stanbon(75887)

how do i write the linear equations in
slope -intercept form for 16=0.5x-2y-2
--------------------
Solve for "y":
2y = 0.5x-2-16
2y = 0.5x-18
y = (1/4)x - 9
--------------
Cheers,
Stan H.
===========

Answer by oberobic(2304)

16 = .5x -2y -2
.
16 = 1/2*x -2y -2
.
2y + 16 = 1/2*x -2
.
2y = 1/2*x -18
.
y = 1/4*x - 9
.